You own 1,100 shares of stock in Avondale Corporation. You will receive a dividend of $1.70 per share in one year. In two years, Avondale will pay a liquidating dividend of $75 per share. The required return on Avondale stock is 20 percent.

a) Ignoring taxes, what is the current share price of your stock? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

Share price $ ???

b) If you would rather have equal dividends in each of the next two years, how many shares would you sell in one year? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

Number of shares: ???

c) What would your cash flow be for each year for the next two years if you create equal homemade dividends? ???

Hint: Dividends will be in the form of an annuity. (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Cash flow $ ReferenceseBook & Resources Worksheet
Answer #1
a] Current share price = 1.70/1.2+75/1.2^2 = $              53.50
b] Total dividend receivable in 1 year = 1100*1.7 = $         1,870.00
Yearly cash flow required [form c below] $       38,520.00
Balance to be received from sale of shares $       36,650.00
Price per share at end of year 1 = 75/1.2 = $              62.50
Number of shares to be sold = 36650/62.50 = 586.40 Shares
c] Current price of 1100 shares = 1100*53.50 = $       58,850.00
Yearly cash flow = 58850*0.20*1.2^2/(1.2^2-1) = $       38,520.00